if f(1)= 3 and f '(1)=-2 find d/dx [x^2 f(x) ] when x=1


1
Expert's answer
2021-01-07T16:05:26-0500

Let "g(x)=x^2f(x)"


Differentiate "g(x)=x^2f(x)" with respect to "x" using product rule as,


"g'(x)=\frac{d}{dx}g(x)"


"=\frac{d}{dx}(x^2f(x))"


"=x^2\frac{d}{dx}(f(x))+f(x)\frac{d}{dx}(x^2)"


Use "\frac{d}{dx}(x^n)=nx^{n-1}" to obtain,


"=x^2f'(x)+f(x)(2x)"


"=x^2f'(x)+2xf(x)"


Derivative at "x=1" is,


"g'(1)=(1^2)f'(1)+2(1)f(1)"


"=f'(1)+2f(1)"


Plug "f'(1)=-2" and "f(1)=3" to obtain "g'(1)" as,


"g'(1)=-2+2(3)"


"=-2+6"


"=4"


Therefore, the derivative of "g(x)=x^2f(x)" at "x=1" is "4" .